OpenStudy (anonymous):

ulio has started his very own floral business. He sells centerpieces for $19.50 each. His fixed costs are $500 per month, and each centerpiece cost $8 to produce. Determine your break-even point. Suppose you wish to break even by selling only 25 centerpieces. At what price would you need to sell each centerpiece? a. Julio must sell 25 centerpieces at $35 each to break even. b. Julio must sell 25 centerpieces at $25 each to break even. c. Julio must sell 25 centerpieces at $40 each to break even. d. Julio must sell 25 centerpieces at $28 each to break even.

OpenStudy (igreen):

His break point is $500. He sells 25 centerpieces. 500 / 25 = ? @yungkrystal

OpenStudy (anonymous):

20?

OpenStudy (igreen):

Yes, but it costs $8 to make each one..so..

OpenStudy (igreen):

If he sold them for $25 each, then he would be making $17 per centerpiece sold.

OpenStudy (igreen):

17 * 25 = 425, so 25 can't be our answer.

OpenStudy (igreen):

If he sold them for $28, he would be making $20 per centerpiece sold. 20 * 25 = 500. That's even.

OpenStudy (igreen):

@myininaya @dan815 @iambatman Hmm..I'm not so sure..would it be 28 or 35? 28 is just even..

OpenStudy (anonymous):

ok